OpenStudy (anonymous):

A pineapple was found to have two sets of opposing spirals formed by its leaves. The clockwise spiral has x leaves. The number of leaves in the counterclockwise direction is 5 more than that in the clockwise direction. These leaf numbers are consecutive numbers in the Fibonacci sequence. Which equation is true for the number of leaves in the two spirals?

OpenStudy (anonymous):

x+5/x=1.6 5/x=1.6 5+5/5=1.6 x/5=1.6

OpenStudy (anonymous):

@cwrw238 @brookester6

OpenStudy (anonymous):

C)

OpenStudy (anonymous):

Omg Thank You! I Hate Math.

OpenStudy (anonymous):

lol yeah it sucks

OpenStudy (anonymous):

It Was Wrong. Ugh. I Dont Understand This.